# url]https://person.dibris.unige.it/ricca-filippo/

### v1

@@ Abstract
Objectives: This study surveys grey literature to explore how AI is adopted in TA, focusing on the problems it solves, its solutions, and the available tools. Additionally, the study gather expert insights to understand AI’s current and future role in TA.
Methods: We reviewed over 3,600 grey literature sources over five years, including blogs, white papers, and user manuals, and finally filtered 342 documents to develop taxonomies of TA problems and AI solutions. We also cataloged 100 AI-driven TA tools and interviewed five expert software testers to gain insights into AI’s current and future role in TA.
@@ 1 Introduction
The integration of Artificial Intelligence (AI) and Machine Learning (ML) into TA is getting significant attention from researchers and practitioners, recognizing AI’s potential to bridge the gap between human and machine-assisted testing activities [15]. In this paper, we refer to these techniques as Artificial Intelligence assisted Test Automation (AIaTA). AI holds the promise of transforming TA by simplifying or automating various testing activities, including test planning, authoring, development, and maintenance. Despite the growing adoption of AIaTA by companies [47], there remains a limited understanding of the challenges it addresses, the solutions it offers, and the existing tools and their integration with the software development process.
Several secondary studies have reviewed existing work on AIaTA [18, 21, 30, 39, 56]. This paper is a secondary study that specifically focuses on the grey literature [11] to capture practitioners’ perspectives on the adoption of AIaTA. The main goal of our work is to understand the available AIaTA tools, the problems they address, and the innovative solutions they offer. Therefore, we conducted a survey of the grey literature to collect, consolidate, and organize existing AI practices for TA. This paper builds upon the grey literature reviews by Ricca et al. [47, 48] by incorporating more recent sources, thus presenting an extended, multi-year grey literature review. This comprehensive review spans several years, demonstrating the generalizability of the results and examining the underlying relationships among TA problems, AI-based solutions, and existing tools. To gather further insights on the application of AIaTA, we conducted interviews with five researchers and practitioners to collect their experience. These interviews were recorded, transcribed, and analyzed to identify the key problems, solutions, and tools discussed by the interviewees. Thus, the main contribution of this extended paper is a set of taxonomies of problems, solutions, and tools regarding AIaTA, which are validated by experts. To our knowledge, this is the first work that includes interviews with developers on the usage of AI related to E2E testing of web apps.
@@ 3 Empirical Study
RQ4 (Generalizability). How generalizable are the taxonomies related to issues, solutions, and tools constructed to address the previous RQs?
RQ5 (Problems vs. Solutions). What is the relationship between TA problems and solutions in AIaTA?
RQ6 (Tools vs. Solutions). How are AIaTA tools and solutions interconnected?
@@
RQ4 assesses the degree of generalizability of the built taxonomies by validating them on a set of studies that were not used during its development.
RQ5 maps TA problems and issues with proposed AI-based solutions. To answer RQ5, we mapped the problems and issues identified in RQ1 with their proposed AI-based solution identified in RQ2.
Finally, the last research question RQ6 maps such proposed AI-based solutions towards existing tools that implement them. To answer RQ6, we mapped the proposed AI-based solution identified in RQ2 with the implemented tools supporting them, as listed in RQ3.
@@ 3.1 Procedure
To review the grey literature relevant to our research questions, we iterated the five-step procedure defined in our initial publication [47] three times, with an approximately one-year interval between each iteration, and further refined it in our extended study [48]. Figure 1 shows these five steps:
@@ 3.1.1 Google Search
The first part of the query string is characterized by words related to artificial intelligence and machine learning, whereas the second part is related to automated testing. All relevant documents contained an instance of each keyword from each part of the string (AND operator), whereas keywords within the same part were ORed. For each query, the first 15 pages of results were scraped, each having 10 documents. We conducted eight queries, which accounted for 1,200 documents that were analyzed overall (150 documents for each query). The output of this step is a list of candidate web documents potentially regarding TA conducted with AI-based solutions. Table 1 (column “# Date”) reports the date on which the search was conducted.
@@ 3.1.2 Document Selection
The list of candidate web documents has been analyzed according to a set of inclusion and exclusion criteria, for filtering out irrelevant documents. We mainly considered three inclusion criteria: (i) the document needs to investigate AI or ML tools or methods that can support TA; (ii) the document should apply to either capture-replay, programmable (or script-based), visual, or combinations of these testing approaches; (iii) tools’ websites and presentations are included as long as they specify useful information, and (iv) presentations and slide decks are included only if in scope. We mainly excluded documents (exclusion criteria) if: (i) they are scientific peer-reviewed papers; (ii) they are not written in the English language; (iii) they provide guidelines for using AI within manual testing; (iv) they are videos or books; (v) they are part of the website that requires registration; and (vi) they contain only generic information on AI or software testing, without relating the two concepts. These criteria have been applied by analyzing each document obtained in the initial search on Google, thus only those documents that provide direct evidence about the study’s objective were retained. The output of this step is a list of relevant candidate documents regarding TA conducted with AI-based solutions. Table 1 (column “# Analyzed”) reports the number of documents selected and then analyzed in this step for each iteration. In parentheses, we have reported the cumulative total value across the various iterations.
@@
Table 1: Summary of data related to the taxonomies.
Iteration Date # Analyzed # Problems # Solutions # Tools
First [47] Sept. 2020 156 35 35 50
Second [48] Febr. 2023 95 (251) 5 (40) 9 (44) 17 (67)
Third April 2024 91 (342) 7 (47) 15 (59) 33 (100)
3.1.3 Data Extraction
In this step, each relevant candidate document has been read and analyzed in detail and the following information has been collected: (1) web link to the document, (2) name and surname, if any, about the author(s) of the document, (3) the publication date of the document (if available), (4) the type of the document (e.g., blog post, interview transcript), (5) the test automation tool(s) investigated or described in the document, (6) the testing level discussed (e.g., unit, integration, system, acceptance), (7) the test automation problems and issues addressed, and (8) the solution offered. For each iteration, the authors conducted a pilot study by labeling a randomly selected sample of 10 documents. The consensus on the labels was high in all the cases, allowing the authors to proceed with the analysis independently on separate sets of documents for the subsequent phases. In detail, we build a tabular representation of the data extracted from the documents implemented as an online spreadsheet on Google Docs. Figure 2 shows an image of the online spreadsheet on Google Docs we used to gather data, which details the contribution aspects. Each row of the table reports the information collected for a document and an individual problem or solution. Hence, more rows can be used to detail several TA problem-solution tuples found in a single document. During the mapping process, the authors reused existing labels whenever applicable to avoid introducing nearly identical labels for the same TA problem/solution and to maintain consistent naming conventions. We recall that the set of TA problems and TA solutions considered has been defined iteratively by starting from a small set of documents (as documented in [47]). Each newly identified problem and/or solution was incrementally considered, and the document was re-analyzed if needed.
@@ 3.1.4 Taxonomy Creation
By collecting the identified tuples of test automation problems and solutions, as well as the list of existing tools, two taxonomies have been constructed to answer RQ1 and RQ2, following a systematic process [14], while the list of tools has been used to answer RQ3. To build the taxonomies, we clustered related TA problems and/or TA solutions, thus identifying categories of problems and solutions. Then, parent categories have been created, by following a specialization relationship between categories and subcategories. Especially, at the beginning of the process, some iterations between this step and the previous one, related to Data extraction, were needed to clearly identify the relevant information. The remaining columns of Table 1 report, for each iteration, the number of identified categories of problems (column “# Problems”) and solutions (“# Solutions”) in the two taxonomies and the number of identified tools (“# Tools”). In parentheses, we have reported the cumulative total value across the various iterations.
@@ 3.1.5 Data Analysis
We answered the first three research questions (RQ1, RQ2 and RQ3) by analyzing the two taxonomies of problems and solutions and the list of tools. In order to answer RQ4, about the generalizability of the built taxonomies, we analyzed the set of categories, sub-categories, and values added in each of the three iterations. Each row in Table 1 reports the data concerning a specific iteration of the experiment. Further analysis has been conducted after the construction of the taxonomies by considering the information related to the tuples of TA problems, solutions, and tools. In particular, several Sankey diagrams [50] have been constructed aiming at visualizing the possible interconnections between AI-based solutions and TA problems (to answer RQ5), and AI-based tools and AI-based solutions (to answer RQ6). In a Sankey diagram, the relations between nodes are shown by links that connect input and output nodes (in our case, respectively, solutions–problems and tools–solutions), while the width of a link indicates the relevance or magnitude of the relationship. By using the Sankey diagrams, we identified trends and patterns in the relations among TA solutions, problems, and tools, thus being able to answer RQ5 and RQ6. More specifically, we employed pandas data-frames [34] to filter information related to the problems addressed and solutions provided (RQ5), as well as the tools used (RQ6). We excluded entries where the problem, solution, or tool was either unspecified or too generic. For the relevant entries, we counted each pairwise combination of <problem, solution> and <tool, solution>, treating each instance of a problem-solution pair or tool-solution pair as a unique connection. We utilized the plotly library [17] to generate the Sankey diagrams [50], though it is important to note that the diagrams presented in this paper focus only on the most prominent connections for the sake of space and readability, while the complete set of diagrams encompasses all collected sources.
@@ 4.1 RQ1 (Issues/problems)
Overall, in the three iterations, we identified 917 individual occurrences into the seven main categories: (1) test planning, 5.6% of occurrences; (2) test design, 5.9% of occurrences; (3) test authoring, 24.1% of occurrences; (4) test type, 0.3% of occurrences; (5) test execution, 12.8% of occurrences; (6) test closure, 10.4% of occurrences; and (7) test maintenance, 20% of occurrences. Other two categories such as Unspecified and Generic collected respectively 14.5% and 5.9% of occurrences.
@@
Untested code and Manual data creation are key aspects to optimize the generation of effective test cases. On the one side, in fact, it is impossible to test everything in a software application, so strategies and techniques to identify the less tested portion of the application are fundamental for reducing the overall effort by focusing on where it is needed. For instance, it has been demonstrated that test suites with low coverage of the app code have a lower chance of detecting bugs [5]. On the other side, the use of high-quality test data is a critical part of testing [49], but producing such high-quality test data can be time-consuming.
@@ 4.2 RQ2 (Solutions/approaches)
The solutions have been grouped into eight main categories. In the table, TA solutions and their occurrences are listed for iterations. Overall, in the three iterations, we identified 972 individual occurrences into the eight main categories: (1) test generation, 31% of occurrences; (2) test oracles, 7.2% of occurrences; (3) debugging, 15.3% of occurrences; (4) test maintenance, 20.6% of occurrences; (5) test process, 2% of occurrences; (6) test execution, 4.1% of occurrences; (7) test optimization, 1.2% of occurrences; and (8) test type, 0.4% of occurrences. Other two categories such as Unspecified and Generic collected respectively 12.2% and 5.6% of occurrences.
@@
Another representative subcategory, even if less numerous, is related to test maintenance, i.e., Self-healing mechanisms (6.4%). Self-healing refers to the capability of automatically applying corrective actions when a test script is broken, for instance, after an application maintenance or evolution task, without human intervention. In fact, the continuous evolution of modern applications breaks test scripts, thus requiring a large effort to repair such test scripts. Following the analysis of the documents, we further split this subcategory into two distinct categories: Self-healing test script (5.6%) and Smart Locators (2.4%). The adoption of smart locators can be relevant to prevent the need for repairing interventions. By using multiple constructs that can be updated dynamically as the application evolves, in fact, smart locators tend to be resilient to test breakages. In detail, by using multiple attributes per web element to locate, smart locators improve the robustness of test scripts.

### v2

@@ Abstract
Objectives: Given the prevalent usage of AI in industry, sources of truth are held in grey literature as well as the minds of professionals, stakeholders, developers, and end-users. To this aim, our study surveys grey literature to explore how AI is adopted in TA, focusing on the problems it solves, its solutions, and the available tools. Additionally, the study is complemented by expert insights.
Methods: Over five years, we reviewed over 3,600 grey literature sources, including blogs, white papers, and user manuals, and finally filtered 342 documents to develop taxonomies of TA problems and AI solutions. We also cataloged 100 AI-driven TA tools and interviewed five expert software testers to gain insights into AI’s current and future role in TA.
@@ 1 Introduction
The integration of Artificial Intelligence (AI) and Machine Learning (ML) into TA is getting significant attention from researchers and practitioners, recognizing AI’s potential to bridge the gap between human and machine-assisted testing activities [20, 58]. In this paper, we refer to these techniques as Artificial Intelligence assisted Test Automation (AIaTA). AI holds the promise of transforming TA by simplifying or automating various testing activities, including test planning, authoring, development, and maintenance. Despite the growing adoption of AIaTA by companies [55], there remains a limited understanding of the challenges it addresses, the solutions it offers, and the existing tools and their integration with the software development process.
Several secondary studies have reviewed existing work on AIaTA [26, 29, 38, 47, 67]. This paper is a secondary study that focuses on the grey literature [16] to capture practitioners’ perspectives on the adoption of AIaTA. The main goal of our work is to understand the available AIaTA tools, the problems they address, and the innovative solutions they offer, from the developers’ perspective.
In our previous work [56, 55], of which this article is an extension, we surveyed the grey literature to collect, consolidate, and organize existing AI practices for TA. This paper builds upon the grey literature reviews by Ricca et al. [55, 56] by incorporating more recent sources, thus presenting an extended, multi-year grey literature review. We provide details on the differences between the prior papers and this article. First, this article contains an expanded empirical evaluation (Section 3) as we extended the study to sources from the year 2024. This comprehensive review spans several years, demonstrating the generalizability of the results and examining the underlying relationships among TA problems, AI-based solutions, and existing tools. Second, to gather further insights on the application of AIaTA, we conducted interviews with five researchers and practitioners to collect their experience. These interviews were recorded, transcribed, and analyzed to identify the key problems, solutions, and tools discussed by the interviewees. Thus, the main contribution of this extended paper is a set of taxonomies of problems, solutions, and tools regarding AIaTA, which are corroborated by experts. To our knowledge, this is the first work that includes interviews with developers on the usage of AI for TA.
@@ 3 Empirical Study
RQ4 (Problems vs. Solutions). What is the relationship between TA problems and solutions in AIaTA?
RQ5 (Tools vs. Solutions). How are AIaTA tools and solutions interconnected?
@@
RQ4 maps TA problems and issues with proposed AI-based solutions. To answer RQ5, we mapped the problems and issues identified in RQ1 with their proposed AI-based solution identified in RQ2.
Finally, the last research question RQ5 maps such proposed AI-based solutions towards existing tools that implement them. To answer RQ6, we mapped the proposed AI-based solution identified in RQ2 with the implemented tools supporting them, as listed in RQ3.
@@ 3.1 Procedure
To review the grey literature relevant to our research questions, we carried out a selection procedure designed according to the guidelines by Garousi et al. [16]. Particularly, we iterated the five-step procedure defined in our initial publication [55] three times, with an approximately one-year interval between each iteration, and further refined it in our extended study [56]. Figure 1 shows these five steps:
@@ 3.1.1 Google Search
The first part of the query string is characterized by words related to artificial intelligence and machine learning, whereas the second part is related to automated testing. All relevant documents contained an instance of each keyword from each part of the string (AND operator), whereas keywords within the same part were ORed. Hence, we have applied eight queries overall in each of the three repetitions of the search. For each query, the first 15 pages of results were scraped, each having 10 documents. No relevant documents were found after the 15th page. The eight applied queries led us to collect 1,200 documents for each repetition (150 documents for each query), thus resulting in a total of 3,600 documents collected and then analyzed in the three repetitions. The output of this step is a list of candidate web documents potentially regarding TA conducted with AI-based solutions. Table 1, among other aspects, reports the number of documents collected in this step for each iteration (column “# Docs”) and the date on which the search was conducted (column “# Date”).
@@ 3.1.2 Document Selection
The list of candidate web documents has been analyzed according to a set of inclusion and exclusion criteria, for filtering out irrelevant documents. We mainly considered three inclusion criteria: (i) the document needs to investigate AI or ML tools or methods that can support TA; (ii) the document should apply to either capture-replay, programmable (or script-based), visual, or combinations of these testing approaches; (iii) tools’ websites and presentations are included as long as they specify useful information, and (iv) presentations and slide decks are included only if in scope. We mainly excluded documents (exclusion criteria) if: (i) they are scientific peer-reviewed papers; (ii) they are not written in the English language; (iii) they provide guidelines for using AI within manual testing; (iv) they are videos or books; (v) they are part of the website that requires registration; and (vi) they contain only generic information on AI or software testing, without relating the two concepts. These criteria have been applied by analyzing each document obtained in the initial search with Google, thus only those documents that provide direct evidence about the study’s objective were retained. The output of this step is the list of relevant selected documents regarding TA conducted with AI-based solutions. Table 1 (column “# Selected”) reports the number of documents selected for the analysis in this step for each iteration, among the ones collected in the Google search step. In parentheses, we have reported the cumulative total value across the various iterations. We can see that, in each repetition, we selected around 10% of the collected web documents, for an overall selection of about 9.5% (i.e., 342 out of 3,600 documents) of documents identified in the Google Search phase.
Table 1: Summary of data related to the document search and taxonomies construction.
Iteration Date #Docs #Selected #Non-Con. #Problems #Solutions #Tools
First [55] Sept. 2020 1,200 156 14 35 35 50
Second [56] Febr. 2023 1,200 95 (251) 4 5 (40) 9 (44) 17 (67)
Third Apr. 2024 1,200 91 (342) 6 7 (47) 15 (59) 33 (100)
3.1.3 Data Extraction
In this step, each relevant candidate document has been read and analyzed in detail and the following information has been collected: (1) web link to the document, (2) name and surname, if any, about the author(s) of the document, (3) the publication date of the document (if available), (4) the type of the document (e.g., blog post, interview transcript), (5) the test automation tool(s) investigated or described in the document, (6) the testing level discussed (e.g., unit, integration, system, acceptance), (7) the test automation problems and issues addressed, and (8) the solution offered. Table 2 reports a summary of the most frequent types of documents among the selected ones, e.g., blog posts, and interview transcripts (column “Document Type” in the top part of the table). We can see that most of the documents were web articles or blogs. Overall, we were able to classify most of the document, i.e., 76.9%. Table 2 reports also the number of documents in which the author’s name was not specified (row “Without author(s)” in the bottom part of the table). We can see that it was impossible to identify the name of the authors for a non trivial number of documents, overall for 37.1% of documents, with an increased trend in the last repetition.
For each iteration, the authors conducted a pilot study by labeling a randomly selected sample of 10 documents. Each author labeled independently the sample of documents (e.g., type of TA problems and TA solutions), then during a meeting each document that was labeled differently by at least two authors was re-analyzed by all the authors together and the labels were discussed for reaching an agreement. These discussions were also used to refine the descriptions of the different categories of the taxonomies, thus reducing ambiguous definitions. Only in a few of the compared cases (4.5%), we started the discussion for reaching an agreement since we observed different labels. Instead in the other cases (95.5%), we observed the same labels for at least two classifiers, out of three, and for all three classifiers in the majority of the cases (59.1%). Generally, hence, we observed that the consensus on the classification of the documents was high and the discussions quite limited, allowing the authors to proceed with the analysis independently on separate sets of documents for the subsequent phases.
Table 2: Summary of types of documents.
Document Type # First [55] # Second [56] # Third
Blog 38 24 35
Interview transcript 0 2 0
Manual 0 2 0
Post 2 7 8
Slide 8 0 0
StackOverflow 2 0 0
Technical feed 0 0 3
Tool 6 3 1
Web article 46 40 30
White-paper 5 0 0
Total 108 (69.2%) 78 (82.1%) 77 (84.6%)
263 (76.9%)
Without author(s) 53 (33.9%) 22 (23.1%) 52 (57.1%)
127 (37.1%)
@@ 3.1.3 Data Extraction
In detail, we build a tabular representation of the data extracted from the documents implemented as an online spreadsheet on Google Docs. Figure 2 shows an image of the online spreadsheet on Google Docs we used to gather data, which details the contribution aspects. Each row of the table reports the information collected for a document and an individual problem or solution. Hence, more rows can be used to detail several TA problem-solution tuples found in a single document. During the mapping process, the authors reused existing labels whenever applicable to avoid introducing nearly identical labels for the same TA problem/solution and to maintain consistent naming conventions. We recall that the set of TA problems and TA solutions considered has been defined iteratively by starting from a small set of documents (as documented in [55]). Each newly identified problem and/or solution was incrementally considered, and the document was re-analyzed if needed.
In the three iterations, not all documents contributed to the identification of relevant TA problems, solutions, and tools supporting them. Table 1 (column “# Non-Con.”) reports the number of documents that did not contribute to the taxonomies in each iteration. We can notice that the reported numbers are rather limited and, in particular, they decrease between the first iteration and the subsequent ones, meaning that the selected documents tend to discuss more and more interesting content.
Table 3: Summary of Credibility and Content Quality.
First [55] Second [56] Third Total
Credibility
Low 39 (42.8%) 39 (57.3%) 36 (42.8%) 12 (4.5%)
Medium 33 (36.2%) 22 (32.2%) 33 (39.2%) 104 (39.8%)
High 19 (20.8%) 7 (10.2%) 15 (17.8.7%) 145 (55.5%)
Content Quality
Low 8 (8.1%) 2 (4.1%) 1 (1.1%) 114 (46.9%)
Medium 32 (32.6%) 35 (47.9%) 37 (41.1%) 88 (36.2%)
High 58 (59.1%) 35 (47.9%) 52 (57.7%) 41 (16.8%)
It is worth noting that the quality of documents typically found in grey literature, such as those analyzed in this study, tends to be more diverse and often more challenging to evaluate compared to traditional systematic literature reviews. It is also more complex to conduct a sort of quality assessment of the used sources. By inspired the quality assessment checklist reported by [16], we focused on: (i) the presence of authors; (ii) the quality and understandability of the content of the document; and (iii) a subjective assessment of the credibility of the document as perceived by the classifiers, i.e., the authors in our case. For each iteration, we collected this information for each selected document. In particular, the content of the document and the document’s credibility are scored in terms of Low/Medium/High according to the judgment of the classifiers. We recall that Table 2 (row “Without author(s)” in the bottom part of the table) reports the number of documents for which an author was not identified.
Table 3 summarizes the results in terms of content quality and document credibility as assessed by the authors. We can notice that overall only a limited amount of documents have been classified as Low in terms of credibility (i.e., 4.5%). However, the classifiers did not provide any judgment for approximately 23% of the documents (i.e., 342 selected documents minus 12+104+145), suggesting uncertainty or low credibility for these sources. Concerning the document content quality, we can notice that overall a large set of documents have been classified as of Low quality (i.e., 46.9%). Moreover, the classifiers did not express any judgment for about 28% (i.e., 342 selected documents minus 114+88+41) of the documents, for which we can assume a doubt to low content quality. These numbers suggest that, on one hand, the overall quality of the considered sources is sufficient for analysis; on the other hand, they confirm that grey literature is not always reliable [16].
@@ 3.1.4 Taxonomy Creation
By collecting the identified tuples of test automation problems and solutions, as well as the list of existing tools, two taxonomies have been constructed to answer RQ1 and RQ2, following a systematic process [19], while the list of tools has been used to answer RQ3. To build the taxonomies, we clustered related TA problems and/or TA solutions, thus identifying categories of problems and solutions. Then, parent categories have been created, by following a specialization relationship between categories and subcategories. Especially, at the beginning of the process, some iterations between this step and the previous one, related to Data extraction, were needed to clearly identify the relevant information. The last columns of Table 1 report, for each iteration, the number of identified categories of problems (column “# Problems”) and solutions (“# Solutions”) in the two taxonomies and the number of identified tools (“# Tools”). In parentheses, we have reported the cumulative total value across the various iterations.
@@ 3.1.5 Data Analysis
We answered the first three research questions (RQ1, RQ2 and RQ3) by analyzing the two taxonomies of problems and solutions and the list of tools. Further analysis has been conducted after the construction of the taxonomies by considering the information related to the tuples of TA problems, solutions, and tools. In particular, several Sankey diagrams [60] have been constructed aiming at visualizing the possible interconnections between AI-based solutions and TA problems (to answer RQ4), and AI-based tools and AI-based solutions (to answer RQ5). In a Sankey diagram, the relations between nodes are shown by links that connect input and output nodes (in our case, respectively, solutions–problems and tools–solutions), while the width of a link indicates the relevance or magnitude of the relationship. By using the Sankey diagrams, we identified trends and patterns in the relations among TA solutions, problems, and tools, thus being able to answer RQ4 and RQ5. More specifically, we employed pandas data-frames [42] to filter information related to the problems addressed and solutions provided (RQ4), as well as the tools used (RQ5). We excluded entries where the problem, solution, or tool was either unspecified or too generic. For the relevant entries, we counted each pairwise combination of <problem, solution> and <tool, solution>, treating each instance of a problem-solution pair or tool-solution pair as a unique connection. We utilized the plotly library [23] to generate the Sankey diagrams [60], though it is important to note that the diagrams presented in this paper focus only on the most prominent connections for the sake of space and readability, while the complete set of diagrams encompasses all collected sources.
@@ 4.1 RQ1 (Issues/problems)
Overall, in the three iterations, we identified 917 individual occurrences (545 after the second iteration, and 423 after the first one) into the seven main categories: (1) test planning, 5.6% of occurrences; (2) test design, 5.9% of occurrences; (3) test authoring, 24.1% of occurrences; (4) test type, 0.3% of occurrences; (5) test execution, 12.8% of occurrences; (6) test closure, 10.4% of occurrences; and (7) test maintenance, 20% of occurrences. Other two categories such as Unspecified and Generic collected respectively 14.5% and 5.9% of occurrences.
@@
Untested code and Manual data creation are key aspects to optimize the generation of effective test cases. On the one side, it is impossible to test everything in a software application, so strategies and techniques to identify the less tested portion of the application are fundamental for reducing the overall effort by focusing on where it is needed. For instance, it has been demonstrated that test suites with low coverage of the app code have a lower chance of detecting bugs [7]. On the other side, the use of high-quality test data is a critical part of testing [57], but producing such high-quality test data can be time-consuming.
@@ 4.2 RQ2 (Solutions/approaches)
The solutions have been grouped into eight main categories. In the table, TA solutions and their occurrences are listed for iterations. Overall, in the three iterations, we identified 972 individual occurrences (607 after the second iteration, and 466 after the first one) into the eight main categories: (1) test generation, 31% of occurrences; (2) test oracles, 7.2% of occurrences; (3) debugging, 15.3% of occurrences; (4) test maintenance, 20.6% of occurrences; (5) test process, 2% of occurrences; (6) test execution, 4.1% of occurrences; (7) test optimization, 1.2% of occurrences; and (8) test type, 0.4% of occurrences. Other two categories such as Unspecified and Generic collected respectively 12.2% and 5.6% of occurrences.
@@
Another representative subcategory, even if less numerous, is related to test maintenance, i.e., Self-healing mechanisms (6.4%). Self-healing refers to the capability of automatically applying corrective actions when a test script is broken, for instance, after an application maintenance or evolution task, without human intervention. The continuous evolution of modern applications breaks test scripts, thus requiring a large effort to repair such test scripts. Following the analysis of the documents, we further split this subcategory into two distinct categories: Self-healing test script (5.6%) and Smart Locators (2.4%). The adoption of smart locators can be relevant to prevent the need for repairing interventions. By using multiple constructs that can be updated dynamically as the application evolves smart locators tend to be resilient to test breakages. In detail, by using multiple attributes per web element to locate, smart locators improve the robustness of test scripts.